OpenStudy (anonymous):

Solve and check 1/2m=7/8 help for my semester exam will give medals

OpenStudy (anonymous):

@hero

OpenStudy (anonymous):

@LaloHao

OpenStudy (anonymous):

x=7/4 or 1 3/4

OpenStudy (anonymous):

Can you show me how you did that

OpenStudy (anonymous):

\[\frac{1}{2}m=\frac{7}{8 }\] Multiply both sides by two \[\frac{1*2}{2}m=\frac{7*2}{8 }\] \[m=\frac{14}{8 }\] Then just simplify \[m=\frac{14}{8}=\frac{7}{4}=1\frac{3}{4}\]

OpenStudy (anonymous):

Thank you